# Weak compactness cardinals for strong logics and subtlety properties of the class of ordinals

### v1

@@ Abstract.
The author would like to thank Will Boney and Victoria Gitman for discussions that motivated much of the content of this paper. In addition, he would like to thank Toshimichi Usuba for helpful comments on earlier versions of the presented results. The author gratefully acknowledges supported by the Deutsche Forschungsgemeinschaft (Project number 522490605).
@@ 1. Introduction
The theory of strong axioms of infinity and the study of extensions of first-order logic are deeply connected through results showing that the ability to generalize fundamental structural features of first-order logic to stronger logics is equivalent to the existence of large cardinals. [^1] Throughout this paper, we refer to properties of cardinal numbers that imply weak inaccessibility as large cardinal properties. A classical example of such a result is a theorem of Magidor in [14] that shows that the existence of a strong compactness cardinal for second-order logic $\mathcal{L}^{2}$ (i.e., the existence of a cardinal $\kappa$ with the property that every unsatisfiable $\mathcal{L}^{2}$-theory contains an unsatisfiable subtheory of cardinality less than $\kappa$) is equivalent to the existence of an extendible cardinal. In [16], Makowsky proved an analog of this result that provides a large cardinal characterization for the existence of strong compactness cardinals for all abstract logics. [^2] The definition of abstract logics and all related model-theoretic notions can be found in Section 5. Recall that Vopěnka’s Principle is the scheme of axioms stating that for every proper class [^3] Throughout this paper, we work in ${\rm{ZFC}}$. Therefore, all classes are definable by first-order formulas in the language of set theory, possibly using sets as parameters. of graphs, there are two distinct members of the class with a homomorphism between them. The validity of this combinatorial principle is known to be characterizable through the existence of certain large cardinals (see [1, Section 4] and [21, Section 6]). Makowsky’s result can then be stated as follows:
@@
The work presented in this paper is motivated by recent results of Boney, Dimopoulos, Gitman and Magidor in [5] that deal with the existence of weak compactness cardinals for abstract logics, i.e., the existence of cardinals $\kappa$ with the property that every unsatisfiable theory of cardinality $\kappa$ contains an unsatisfiable subtheory of cardinality less than $\kappa$. These results show that the existence of weak compactness cardinals for strong logics is connected to the combinatorial concept of subtleness, introduced by Jensen and Kunen in [11]. Remember that an infinite cardinal $\kappa$ is subtle if for every sequence $\langle{A_{\alpha}}~{}|~{}{\alpha<\kappa}\rangle$ with $A_{\alpha}\subseteq\alpha$ for all $\alpha<\kappa$ and every closed unbounded subset $C$ of $\kappa$, there are $\alpha,\beta\in C$ with $\alpha<\beta$ and $A_{\beta}\cap\alpha=A_{\alpha}$. A short argument shows that subtle cardinals are inaccessible limits of totally indescribable cardinals. This large cardinal notion has a canonical class-version, i.e., one defines ${\rm{Ord}}$ is subtle to be the schemes of sentences stating that for every class sequence $\langle{A_{\alpha}}~{}|~{}{\alpha\in{\rm{Ord}}}\rangle$ with $A_{\alpha}\subseteq\alpha$ for all $\alpha\in{\rm{Ord}}$ and every closed unbounded class $C$ of ordinals, there exist $\alpha,\beta\in C$ with $\alpha<\beta$ and $A_{\beta}\cap\alpha=A_{\alpha}$. The results of [5] now show that, in the presence of a definable well-ordering of ${\rm{V}}$, [^4] Note that the existence of such a well-ordering is equivalent to the statement that ${\rm{V}}={\rm{HOD}}_{x}$ for some set $x$ and can therefore be formulated as a single sentence in the language of set theory. the validity of this principle is connected to the existence of weak compactness cardinals for abstract logics:
@@
First, in order to prove results that do not rely on the existence of definable well-orderings, we will make use of the following variation of the above class-principle that was introduce by Bagaria and the author in their work on principles of structural reflection:
@@
Next, we present a characterization of the existence of weak compactness cardinals for all abstract logics in terms of combinatorial properties of the class ${\rm{Ord}}$. This characterization is based on the notion of faintness, introduced by Matet in [17]. Matet defined a cardinal $\kappa$ to be faint if for every sequence $\langle{A_{\alpha}}~{}|~{}{\alpha<\kappa}\rangle$ with $A_{\alpha}\subseteq\alpha$ for all $\alpha<\kappa$ and every $\xi<\kappa$, there are $\xi<\alpha<\beta<\kappa$ with $A_{\beta}\cap\alpha=A_{\alpha}$. The results of [17] then show that a cardinal is faint if and only if it is either subtle or a limit of subtle cardinals. Following earlier patterns, we introduce the following class-version of faintness:
@@
By definition, if ${\rm{Ord}}$ is essentially subtle, then ${\rm{Ord}}$ is essentially faint. Moreover, it is easy to see that the existence of a proper class of subtle cardinals implies that ${\rm{Ord}}$ is essentially faint, because, given a class sequence $\langle{E_{\alpha}}~{}|~{}{\alpha\in{\rm{Ord}}}\rangle$ with $\emptyset\neq E_{\alpha}\subseteq{\mathcal{P}}({\alpha})$ for all $\alpha\in{\rm{Ord}}$ and an ordinal $\xi$, we can pick a subtle cardinal $\kappa>\xi$ and use the Axiom of Choice to find a sequence $\langle{A_{\alpha}}~{}|~{}{\alpha<\kappa}\rangle$ with $A_{\alpha}\in E_{\alpha}$ for all $\alpha<\kappa$. The following result, again proven in Section 6, shows that this principle exactly determines the combinatorial property of ${\rm{Ord}}$ needed for the desired characterization:
@@
As already mentioned earlier, the proofs of the above theorems rely heavily on characterizations of the involved combinatorial class-principles through large cardinal assumptions. This analysis, carried out in Sections 2 and 3 below, can also be used to show that the ability to generalize other structural properties of first-order logic to all abstract logics can also be characterized through the validity of the studied combinatorial principles for ${\rm{Ord}}$. More specifically, motivated by an unpublished result of Stavi that provides an analog to Theorem 1.1 for the existence of Löwenheim–Skolem–Tarski numbers for abstract logics (see Theorem 5.4 below), we will prove analogs of Theorems 1.4 and 1.6 that characterize the existence of strict Löwenheim–Skolem–Tarski numbers, studied by Bagaria and Väänänen in [3], through subtlety properties of ${\rm{Ord}}$ (see Theorem 5.5).
@@
In particular, should it be the case that the axioms of ${\rm{ZFC}}$ prove that there are only set-many subtle cardinals, then these axioms also prove the equivalence of the two principles. Moreover, since the existence of a subtle cardinal clearly implies the existence of a set-sized model of ${\rm{ZFC}}$ in which ${\rm{Ord}}$ is essentially subtle, the above result directly yields the following equiconsistency:
@@
The final topic that will be treated in this paper is the relationship between the existence of weak compactness for abstract logics and the existence of strongly inaccessible cardinals. By combining Theorem 1.1 with the results of [1, Section 4], it is easy to see that the existence of strong compactness cardinals for all abstract logics implies the existence of a proper class of strongly inaccessible cardinals. Moreover, a combination of Theorem 1.4 with results in [2, Section 9] (see Theorem 2.4 below) shows that the same conclusion follows from the assumption that every abstract logic has a stationary class of weak compactness cardinals. In contrast, it turns out that the existence of weak compactness cardinals for abstract logics is compatible with the non-existence of strongly inaccessible cardinals. The corresponding arguments will also allow us to show that there can be incompatible sentences witnessing (2) in Theorem 1.10.
@@
Notice that, in combination with Theorem 1.6, the above theorem shows that the existence of a weak compactness cardinal for second-order logic does not imply the existence of a strongly inaccessible cardinal and this shows that weakly sub-extendible cardinals, introduced by Fuchino and Sakai in their analysis of weak compactness cardinals for $\mathcal{L}^{2}$ in [8], can exist in models of set theory that do not contain strongly inaccessible cardinals.
Finally, while Theorem 1.14 shows that the existence of weak compactness cardinals for all abstract logics does not imply the existence of strongly inaccessible cardinals, our analysis allows us to conclude that the given assumption at least implies the existence of strongly inaccessible cardinals in the inner model ${\rm{HOD}}$:
@@
Note that the proof of this result will actually show that strongly inaccessible cardinals with strong combinatorial properties exist in ${\rm{HOD}}$ (see Corollary 4.5 below).
@@ 2. $C^{(n)}$-strongly unfoldable cardinals
    Given natural numbers $m<n$, every $C^{(n)}$-strongly unfoldable cardinals is $C^{(m)}$-strongly unfoldable.
@@ 3. $C^{(n)}$-Weakly shrewd cardinals
In this section, we will formulate and prove an analog of Theorem 2.4 for the principle ${\rm{Ord}}$ is essentially faint. The starting point of this analysis is the notion of weak shrewdness introduced in [13] as a natural weakening of shrewdness that does not imply the strong inaccessibility of the given cardinal. In [13], a cardinal $\kappa$ is defined to be weakly shrewd if for every formula $\varphi(v_{0},v_{1})$ in the language of set theory, every cardinal $\theta>\kappa$ and every subset $z$ of $\kappa$ with the property that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\theta)$, there exists cardinals $\bar{\kappa}<\bar{\theta}$ such that $\bar{\kappa}<\kappa$ and $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\theta})$. The results of [13, Section 3] then show that every shrewd cardinal is weakly shrewd and every weakly shrewd cardinal is weakly Mahlo. Moreover, these results show that, starting from a subtle cardinal, it is possible to construct a model of set theory that contains a weakly shrewd cardinal that is not strongly inaccessible and therefore also not shrewd. Moreover, it is shown that weakly shrewd cardinals can even exist below the cardinality of the continuum. The results of the paper will show that these consistency results make use of the correct large cardinal hypothesis (see Lemma 4.3 below). Motivated by Lemma 2.3, we strengthen the notion of weak shrewdness in the same way as the transition to $C^{(n)}$-strong unfoldability strengthens the notions of shrewdness.
@@
(1) Let $\kappa$ be a $C^{(n)}$-weakly shrewd cardinal. Fix a formula $\varphi(v_{0},v_{1})$ in the language of set theory, a cardinal $\kappa<\theta\in C^{(m)}$ and $z\subseteq\kappa$ such that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\theta)$. Pick a cardinal $\theta<\vartheta\in C^{(n)}$. Then, in ${\rm{H}}(\vartheta)$, there exists a cardinal $\kappa<\nu\in C^{(m)}$ with the property that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\nu)$. By our assumptions on $\kappa$, we can now find a cardinal $\bar{\vartheta}\in C^{(n)}$ and a cardinal $\bar{\kappa}<\min(\kappa,\bar{\vartheta})$ such that, in ${\rm{H}}(\bar{\vartheta})$, there exists a cardinal $\bar{\kappa}<\bar{\nu}\in C^{(m)}$ with the property that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\nu})$. Since $\bar{\vartheta}\in C^{(n)}$ and $n>0$, we then know that, in ${\rm{V}}$, the ordinal $\bar{\theta}$ is a cardinal in $C^{(m)}$ with the property that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\nu})$.
@@
In combination with [13, Proposition 3.3], the above proposition shows that all weakly $C^{(n)}$-shrewd cardinals are weakly Mahlo. In particular, if a $C^{(n)}$-weakly shrewd cardinal is an element of $C^{(1)}$, then it is strongly inaccessible. Motivated by [13, Lemma 3.1], we now prove an analog of Lemma 2.3 for $C^{(n)}$-weak shrewdness.
@@
By our assumption and the fact that $\mathsf{Sat}$ is uniformly $\Delta_{1}$-definable over models of ${\rm{ZF}}^{-}$, we can now find a cardinal $\bar{\vartheta}\in C^{(n)}$ and a cardinal $\bar{\kappa}<\min(\kappa,\bar{\vartheta})$ with the property that, in ${\rm{H}}(\bar{\vartheta})$, there exists a cardinal $\bar{\kappa}<\bar{\theta}\in C^{(n)}$, an element $\bar{z}$ of ${\rm{H}}(\bar{\theta})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ of cardinality $\bar{\kappa}$ with $\bar{\kappa}\cup\{\bar{\kappa},\bar{z}\}\subseteq X$ and a bijection ${a}:{\bar{\kappa}}\longrightarrow{X}$ with $a(0)=\bar{\kappa}$, $a(\omega\cdot(1+\alpha))=\alpha$ for all $\alpha<\bar{\kappa}$ and the property that $z\cap\bar{\kappa}$ consists of all elements of $\kappa$ of the form ${\prec}{\ell},{\alpha_{0},\ldots,\alpha_{n-1}}{\succ}$ such that $\ell<\omega$ is the Gödel number of a set-theoretic formula with $n$ free variables, $\alpha_{0},\ldots,\alpha_{n-1}<\bar{\kappa}$ and $\mathsf{Sat}(X,\ell,\langle a(\alpha_{0}),\ldots,a(\alpha_{n-1})\rangle)$ holds. Since $\bar{\vartheta}\in C^{(n)}$, we now know that, in ${\rm{V}}$, the ordinal $\bar{\theta}$ is a cardinal in $C^{(n)}$ and the set $X$ is an elementary submodel of ${\rm{H}}(\bar{\theta})$. If we now define
@@
Now, assume that (2) holds, $\varphi(v_{0},v_{1})$ is a formula in the language of set theory, $\kappa<\theta\in C^{(n)}$ is a cardinal and $z$ is a subset of $\kappa$ with the property that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\theta)$. Our assumptions then allows us to find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\bar{\kappa}<\min(\kappa,\bar{\theta})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $z\in{{\rm{ran}}(j)}$. We then know that $z\cap\bar{\kappa}\in X$ and $j(z\cap\bar{\kappa})=z$. But, this shows that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in both $X$ and ${\rm{H}}(\bar{\theta})$. ∎
@@
We close this section by proving the desired characterization of the principle ${\rm{Ord}}$ is essentially faint through the existence of weakly $C^{(n)}$-shrewd cardinals:
@@
First, assume, towards a contradiction, that (1) holds, $n>0$ is a natural number and there exists an ordinal $\xi$ with the property that there are no weakly $C^{(n)}$-shrewd cardinals above $\xi$. Then there is a unique class function $E$ with domain ${\rm{Ord}}$ such that the following statements hold:

### v3

@@ Abstract.
The author would like to thank Will Boney and Victoria Gitman for discussions that motivated much of the content of this paper. In addition, he would like to thank Toshimichi Usuba for helpful comments on earlier versions of the presented results. Finally, the author would like to thank the anonymous referee for a very thorough reading of the manuscript and numerous helpful comments. The author gratefully acknowledges supported by the Deutsche Forschungsgemeinschaft (Project number 522490605).
@@ 1. Introduction
The theory of strong axioms of infinity and the study of extensions of first-order logic are deeply connected through results showing that the ability to generalize fundamental structural features of first-order logic to stronger logics is equivalent to the existence of large cardinals. [^1] Throughout this paper, we refer to properties of cardinal numbers that imply weak inaccessibility as large cardinal properties. A classical example of such a result is a theorem of Magidor in [14] that shows that the existence of a strong compactness cardinal for second-order logic $\mathcal{L}^{2}$ (i.e., the existence of a cardinal $\kappa$ with the property that every unsatisfiable $\mathcal{L}^{2}$-theory contains an unsatisfiable subtheory of cardinality less than $\kappa$) is equivalent to the existence of an extendible cardinal. In [16], Makowsky proved an analog of this result that provides a large cardinal characterization of the existence of strong compactness cardinals for all abstract logics. [^2] The definition of an abstract logic and all related model-theoretic notions can be found in Section 5. Recall that Vopěnka’s Principle is the scheme of axioms stating that for every proper class [^3] Throughout this paper, we work in ${\rm{ZFC}}$. Therefore, all classes are definable by first-order formulas in the language of set theory, possibly using sets as parameters. of graphs, there are two distinct members of the class with a homomorphism between them. The validity of this combinatorial principle is known to be characterizable through the existence of certain large cardinals (see [1, Section 4] and [21, Section 6]). Makowsky’s result now states the following:
@@
The work presented in this paper is motivated by recent results of Boney, Dimopoulos, Gitman and Magidor in [5] that deal with the existence of weak compactness cardinals for abstract logics, i.e., the existence of cardinals $\kappa$ with the property that every unsatisfiable theory of cardinality $\kappa$ contains an unsatisfiable subtheory of cardinality less than $\kappa$. These results show that the existence of weak compactness cardinals for strong logics is connected to the combinatorial concept of subtleness, introduced by Jensen and Kunen in [11]. Remember that an infinite cardinal $\kappa$ is subtle if for every sequence $\langle{A_{\alpha}}~{}|~{}{\alpha<\kappa}\rangle$ with $A_{\alpha}\subseteq\alpha$ for all $\alpha<\kappa$ and every closed unbounded subset $C$ of $\kappa$, there are $\alpha,\beta\in C$ with $\alpha<\beta$ and $A_{\beta}\cap\alpha=A_{\alpha}$. A short argument shows that subtle cardinals are inaccessible limits of totally indescribable cardinals. This large cardinal notion has a canonical class-version, i.e., one defines ${\rm{Ord}}$ is subtle to be the scheme of sentences stating that for every class sequence $\langle{A_{\alpha}}~{}|~{}{\alpha\in{\rm{Ord}}}\rangle$ with $A_{\alpha}\subseteq\alpha$ for all $\alpha\in{\rm{Ord}}$ and every closed unbounded class $C$ of ordinals, there exist $\alpha,\beta\in C$ with $\alpha<\beta$ and $A_{\beta}\cap\alpha=A_{\alpha}$. The results of [5] now show that, in the presence of a definable well-ordering of ${\rm{V}}$, [^4] Note that the existence of such a well-ordering is equivalent to the statement that ${\rm{V}}={\rm{HOD}}_{x}$ for some set $x$ and can therefore be formulated as a single sentence in the language of set theory. the validity of this principle is connected to the existence of weak compactness cardinals for abstract logics:
@@
First, in order to prove results that do not rely on the existence of definable well-orderings, we will make use of the following variation of the above class-principle that was introduced by Bagaria and the author in their work on principles of structural reflection:
@@
Next, we present a result showing that the statement that every abstract logic has at least one weak compactness cardinal is also equivalent to a combinatorial property of the class ${\rm{Ord}}$. This property generalizes the notion of faintness, introduced by Matet in [17], from cardinals to the class ${\rm{Ord}}$. Matet defined a cardinal $\kappa$ to be faint if for every sequence $\langle{A_{\alpha}}~{}|~{}{\alpha<\kappa}\rangle$ with $A_{\alpha}\subseteq\alpha$ for all $\alpha<\kappa$ and every $\xi<\kappa$, there are $\xi<\alpha<\beta<\kappa$ with $A_{\beta}\cap\alpha=A_{\alpha}$. The results of [17] then show that a cardinal is faint if and only if it is either subtle or a limit of subtle cardinals. Following earlier patterns, we introduce the following class-version of faintness:
@@
By definition, if ${\rm{Ord}}$ is essentially subtle, then ${\rm{Ord}}$ is essentially faint. Moreover, it is easy to see that the existence of a proper class of subtle cardinals implies that ${\rm{Ord}}$ is essentially faint: Given a class sequence $\langle{E_{\alpha}}~{}|~{}{\alpha\in{\rm{Ord}}}\rangle$ with $\emptyset\neq E_{\alpha}\subseteq{\mathcal{P}}({\alpha})$ for all $\alpha\in{\rm{Ord}}$ and an ordinal $\xi$, the given assumption allows us to pick a subtle cardinal $\kappa>\xi$ and we can then use the Axiom of Choice to find a sequence $\langle{A_{\alpha}}~{}|~{}{\alpha<\kappa}\rangle$ with $A_{\alpha}\in E_{\alpha}$ for all $\alpha<\kappa$. The subtleness of $\kappa$ then ensures that the closed unbounded subset $(\xi,\kappa)$ of $\kappa$ contains ordinals $\alpha<\beta$ with $A_{\alpha}=A_{\beta}\cap\alpha$. In particular, it follows that there are ordinals $\xi<\alpha<\beta$ and $A\in E_{\beta}$ with $A\cap\alpha\in E_{\alpha}$.
The following result, again proven in Section 6, shows that this principle exactly determines the combinatorial property of ${\rm{Ord}}$ needed for the desired characterization:
@@
As mentioned earlier, the proofs of the above theorems rely heavily on characterizations of the involved combinatorial class-principles through large cardinal assumptions. This analysis, carried out in Sections 2 and 3 below, can also be used to show that the ability to generalize other structural properties of first-order logic to all abstract logics can also be characterized through the validity of the studied combinatorial principles for ${\rm{Ord}}$. More specifically, motivated by an unpublished result of Stavi that provides an analog to Theorem 1.1 for the existence of Löwenheim–Skolem–Tarski numbers for abstract logics (see Theorem 5.4 below), we will prove analogs of Theorems 1.4 and 1.6 that characterize the existence of strict Löwenheim–Skolem–Tarski numbers, studied by Bagaria and Väänänen in [3], through subtlety properties of ${\rm{Ord}}$ (see Theorem 5.5).
@@
In particular, in the unlikely case that the axioms of ${\rm{ZFC}}$ prove that there are only set-many subtle cardinals, the above theorem implies that these axioms also prove that the principle ${\rm{Ord}}$ is essentially subtle is equivalent to the principle ${\rm{Ord}}$ is essentially faint. Our later results will show that the converse of this implication also holds, i.e., if the axioms of ${\rm{ZFC}}$ are consistent with the existence of a proper class of subtle cardinals, then the given principles are not provably equivalent (see Theorem 1.10 below).
Since the existence of a subtle cardinal clearly implies the existence of a set-sized model of ${\rm{ZFC}}$ in which ${\rm{Ord}}$ is essentially subtle, Theorem 1.7 directly yields the following equiconsistency:
@@
The final topic that will be treated in this paper is the relationship between the existence of weak compactness cardinals for abstract logics and the existence of strongly inaccessible cardinals. By combining Theorem 1.1 with the results of [1, Section 4], it is easy to see that the existence of strong compactness cardinals for all abstract logics implies the existence of a proper class of strongly inaccessible cardinals. Moreover, a combination of Theorem 1.4 with results in [2, Section 9] (see Theorem 2.4 below) shows that the same conclusion follows from the assumption that every abstract logic has a stationary class of weak compactness cardinals. In contrast, it turns out that the existence of weak compactness cardinals for abstract logics is compatible with the non-existence of strongly inaccessible cardinals. The corresponding arguments will also allow us to show that there can be incompatible sentences witnessing (2) in Theorem 1.10.
@@
While Theorem 1.14 shows that the existence of weak compactness cardinals for all abstract logics does not imply the existence of strongly inaccessible cardinals, our analysis allows us to conclude that the given assumption at least implies the existence of strongly inaccessible cardinals in the inner model ${\rm{HOD}}$:
@@
Note that the proof of this result will actually show that strongly inaccessible cardinals with certain strong combinatorial properties exist in ${\rm{HOD}}$ (see Corollary 4.5 below).
@@ 2. $C^{(n)}$-strongly unfoldable cardinals
    Given natural numbers $m<n$, every $C^{(n)}$-strongly unfoldable cardinal is $C^{(m)}$-strongly unfoldable.
@@ 3. $C^{(n)}$-Weakly shrewd cardinals
In this section, we will formulate and prove an analog of Theorem 2.4 for the principle ${\rm{Ord}}$ is essentially faint. The starting point of this analysis is the notion of weak shrewdness introduced in [13] as a natural weakening of shrewdness. This large cardinal property still entails many of the interesting combinatorial properties of strongly unfoldable cardinals, but it no longer implies that the given cardinal is strongly inaccessible or an element of $C^{(2)}$. In [13], a cardinal $\kappa$ is defined to be weakly shrewd if for every formula $\varphi(v_{0},v_{1})$ in the language of set theory, every cardinal $\theta>\kappa$ and every subset $z$ of $\kappa$ with the property that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\theta)$, there exist cardinals $\bar{\kappa}<\bar{\theta}$ such that $\bar{\kappa}<\kappa$ and $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\theta})$. The results of [13] show that when we modify this definition to also demand that we can find cardinals $\bar{\theta}$ with the given property below $\kappa$ (as it is the case in the definition of shrewdness), we obtain a strictly stronger large cardinal property. In particular, the canonical argument showing that shrewd cardinal are strongly inaccessible cannot be adapted to weakly shrewd cardinals.
The results of [13, Section 3] show that every shrewd cardinal is weakly shrewd and every weakly shrewd cardinal is weakly Mahlo. Moreover, these results show that, starting from a subtle cardinal, it is possible to construct a model of set theory that contains a weakly shrewd cardinal that is not strongly inaccessible and therefore also not shrewd. Moreover, it is shown that weakly shrewd cardinals can even exist below the cardinality of the continuum. The results of this paper will show that these consistency results make use of the correct large cardinal hypothesis (see Lemma 4.3 below). Motivated by Lemma 2.3, we strengthen the notion of weak shrewdness in the same way as the transition to $C^{(n)}$-strong unfoldability strengthens the notions of shrewdness.
@@
(1) Let $\kappa$ be a $C^{(n)}$-weakly shrewd cardinal. Fix a formula $\varphi(v_{0},v_{1})$ in the language of set theory, a cardinal $\kappa<\theta\in C^{(m)}$ and $z\subseteq\kappa$ such that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\theta)$. Pick a cardinal $\theta<\vartheta\in C^{(n)}$. Then, in ${\rm{H}}(\vartheta)$, there exists a cardinal $\kappa<\nu\in C^{(m)}$ with the property that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\nu)$. By our assumptions on $\kappa$, we can now find a cardinal $\bar{\vartheta}\in C^{(n)}$ and a cardinal $\bar{\kappa}<\min(\kappa,\bar{\vartheta})$ such that, in ${\rm{H}}(\bar{\vartheta})$, there exists a cardinal $\bar{\kappa}<\bar{\nu}\in C^{(m)}$ with the property that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\nu})$. Since $\bar{\vartheta}\in C^{(n)}$ and $n>0$, we then know that, in ${\rm{V}}$, the ordinal $\bar{\nu}$ is a cardinal in $C^{(m)}$ with the property that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in ${\rm{H}}(\bar{\nu})$.
@@
In combination with [13, Proposition 3.3], the above proposition shows that all $C^{(n)}$-weakly shrewd cardinals are weakly Mahlo. In particular, if a $C^{(n)}$-weakly shrewd cardinal is an element of $C^{(1)}$, then it is strongly inaccessible. Motivated by [13, Lemma 3.1], we now prove an analog of Lemma 2.3 for $C^{(n)}$-weak shrewdness.
@@
By our assumption and the fact that $\mathsf{Sat}$ is uniformly $\Delta_{1}$-definable over models of ${\rm{ZF}}^{-}$, we can now find a cardinal $\bar{\vartheta}\in C^{(n)}$ and a cardinal $\bar{\kappa}<\min(\kappa,\bar{\vartheta})$ with the property that, in ${\rm{H}}(\bar{\vartheta})$, there exists a cardinal $\bar{\kappa}<\bar{\theta}\in C^{(n)}$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ of cardinality $\bar{\kappa}$ with $\bar{\kappa}+1\subseteq X$ and a bijection ${a}:{\bar{\kappa}}\longrightarrow{X}$ with $a(0)=\bar{\kappa}$, $a(\omega\cdot(1+\alpha))=\alpha$ for all $\alpha<\bar{\kappa}$ and the property that the set $z\cap\bar{\kappa}$ consists of all elements of $\kappa$ of the form ${\prec}{\ell},{\alpha_{0},\ldots,\alpha_{n-1}}{\succ}$ such that $\ell<\omega$ is the Gödel number of a set-theoretic formula with $n$ free variables, $\alpha_{0},\ldots,\alpha_{n-1}<\bar{\kappa}$ and $\mathsf{Sat}(X,\ell,\langle a(\alpha_{0}),\ldots,a(\alpha_{n-1})\rangle)$ holds. Since $\bar{\vartheta}\in C^{(n)}$, we now know that, in ${\rm{V}}$, the ordinal $\bar{\theta}$ is a cardinal in $C^{(n)}$ and the set $X$ is an elementary submodel of ${\rm{H}}(\bar{\theta})$. If we now define
@@
Now, assume that (2) holds, $\varphi(v_{0},v_{1})$ is a formula in the language of set theory, $\kappa<\theta\in C^{(n)}$ is a cardinal and $z$ is a subset of $\kappa$ with the property that $\varphi(z,\kappa)$ holds in ${\rm{H}}(\theta)$. Our assumptions then allow us to find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\bar{\kappa}<\min(\kappa,\bar{\theta})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $z\in{{\rm{ran}}(j)}$. We then know that $z\cap\bar{\kappa}\in X$ and $j(z\cap\bar{\kappa})=z$. But, this shows that $\varphi(z\cap\bar{\kappa},\bar{\kappa})$ holds in both $X$ and ${\rm{H}}(\bar{\theta})$. ∎
@@
We close this section by proving the desired characterization of the principle ${\rm{Ord}}$ is essentially faint through the existence of $C^{(n)}$-weakly shrewd cardinals:
@@
First, assume, towards a contradiction, that (1) holds, $n>0$ is a natural number and there exists an ordinal $\xi$ with the property that there are no $C^{(n)}$-weakly shrewd cardinals above $\xi$. Then there is a unique class function $E$ with domain ${\rm{Ord}}$ such that the following statements hold: